Seawater of salinity 35,000 ppm at 35oC enters a membrane module of area 1 m2 at a flow rate of 36,000 m3/d. 75% of the flow rate of the seawater feed is retained in the brine reject. If the salt rejection is 99%, calculate:
a) Permeatesalinity
b) Brine salinity
c) Salt passage
d) Recoveryratio
e) Osmoticpressureoffeedseawater
f) Salt flux through membrane
-6 3 2
Salt permeability = 2.03x10 m /m .s
Assume density of seawater is 1000 g/L
Answer #1

Page=2 Salt Content in Feed je SC, - S, XF, = 35,000 x 36,000 = 12661079/day (.1PPM=19/m3) We know that sq=0.99 SC, 1=0.99 (Page-3 * Salt Passage => 2 126x105 1268007 => 0.01 -1% * Recovery Ratio - 9000 36,000 = = = 25% of Osmotic Poessuse offeed SePage- de Osmatic Pressure Os Fx RxT O = 0.598 x 0.082 (35+273) (:R= 0.082 A+). O = 0.598 x 0.082(308) rolk o= 0.998 x 25.256